OpenStudy (kimberly_pr):

Which of the following represents the graph of f(x) = 2x − 5?

OpenStudy (kimberly_pr):

graph of exponential rising up to the right, through the point 0, negative 4 graph of exponential rising up to the right, through the point 0, 6 graph of exponential rising up to the right, through the point 0, 3 graph of exponential rising up to the right, through the point 0, negative 1

OpenStudy (kimberly_pr):

@mathmate @agent0smith

OpenStudy (kimberly_pr):

@ganeshie8

OpenStudy (kimberly_pr):

@hartnn

OpenStudy (kimberly_pr):

@welshfella

OpenStudy (kimberly_pr):

Hey @Photon336 do you know which one of those options represents the function?

OpenStudy (photon336):

f(x) = 2x-5 is a line so try to plug in the points 2(0)-5 = ?

OpenStudy (kimberly_pr):

-5

OpenStudy (photon336):

yeah, this is weird. exponential? the graph is a linear function @ganeshie8

OpenStudy (welshfella):

is it meant to be 2^x + 1 ?

OpenStudy (welshfella):

that would be exponential - it has the variable in the exponent

OpenStudy (welshfella):

2^x - 5 i meant

OpenStudy (kimberly_pr):

Yeah, it's 2^x - 5

OpenStudy (mathmate):

I think the graphs is \(y=2x^{-5}\) Is that right @Kimberly_PR

OpenStudy (kimberly_pr):

No, it's y = 2^x -5

OpenStudy (welshfella):

Hint: all graphs of the form y = a^x rise to the right and pass through the point (0,1) , because a^0 = 1

OpenStudy (mathmate):



OpenStudy (mathmate):

for a graph \(ab^x+k\) there is a vertical shift of k, up if k>0, and down if k<0.

OpenStudy (welshfella):

oh - i forgot to say that a must be > 1

OpenStudy (kimberly_pr):

Ok so since a must be greater than 1 that eliminates options A and D, and thanks for the graph @mathmate

OpenStudy (kimberly_pr):

But now between b and c, which would it be?

OpenStudy (agent0smith):

They give you a point on the graph. Plug in the point into the equation and see if it's true.

OpenStudy (agent0smith):

All you have to do is plug in x=0 and see what y it gives you

OpenStudy (kimberly_pr):

Oh, ok, lol. Makes sense, I'll do that now

OpenStudy (kimberly_pr):

One gives me 7 and the other gives me one, so it would be the one that gives me one right?

OpenStudy (kimberly_pr):

So I think it would be C

OpenStudy (mathmate):

@Kimberly_PR Wait! The graph is for \(y=ab^x\). IF you factor in y=2^x-5, then the whole graph shifts down by 5 units.

OpenStudy (mathmate):

That is a=1, k=-5, what do you get?

OpenStudy (kimberly_pr):

Wait... If it goes through -5 then A would be the answer... but in that graph a isn't greater than 1 @mathmate

OpenStudy (mathmate):

For the graph y=2^x, then a=1, b=2,and the graph passes through (0,1), right?

OpenStudy (kimberly_pr):

The one closest to that is graph C

OpenStudy (mathmate):

I am sorry but I do not confirm or infirm guesses! lol I am asking does (0,1) lie on the graph of y=2^x.

OpenStudy (mathmate):

You can put x=0 into the equation and using 2^0=1 to find out.

OpenStudy (kimberly_pr):

Ok, I will

OpenStudy (kimberly_pr):

I got -4, and the only one that goes through -4 is A

OpenStudy (mathmate):

How do you get -4?

OpenStudy (agent0smith):

y = 2^x - 5 when x=0, y=-4

OpenStudy (kimberly_pr):

^^ exactly

OpenStudy (mathmate):

Right answer, but wrong person! lol

OpenStudy (mathmate):

Yes, @Kimberly_PR you can go ahead and make your choice. There is enough information for you to do so.

OpenStudy (kimberly_pr):

Lol, I just plugged a 0 where the x was in the equation, and that is how I got -4. Thanks

OpenStudy (mathmate):

I know you knew, but @agent0smith beat you to it! lol

OpenStudy (agent0smith):

lol she had it before i did @mathmate

OpenStudy (kimberly_pr):

Yeah, lol. I was about to write it too XD but thanks again to all of you :)

OpenStudy (mathmate):

Sure! we all had a good time at this party! lol